Hook co. has sh.100 million face value of outstanding debt with a coupon of 10% and a par value of sh. 1000. The bonds make annual payments, have a current market of sh. 1025 and are redeemable at par after 10 years. The company also has 1 million shares of common stock with book value per share of $ 35 and a market value per share of $ 50. The current beta of the stock is 1.5 the Treasury bill rate is 5% and the market risk premium is 8%. The company is the 30% tax bracket.

Required

The company's weighted average cost of capital using;

i) book value weights

ii) market value weights
Answer #1

Number of Bonds=100000000/1000=100000 (Face value of the bond/Par value of the bond=100million/1000)

Book Value Market Value
Debt(D) 100000000 (100 Million) 102500000(1025*100000)
Equity(E) 35000000(35*1 million shares) 50000000(50*1 million shares)
Value(V) 135000000 152500000

WACC=(Book value of Equity/value of the firm)*ke+(Book Value of Debt/Value of the firm*Kd*(1-tax rate))

here ke and kd is Cost of equity and cost of debt.

Cost of debt is given 10%=0.1

cost of equity can be calculated by CAPM formula

Ke=Rf+\beta(Rm-Rf)

Rf=treasury Bill rate=0.05, Marker risk premium =Rm=0.08,Beta =1.5

=0.05+1.5(0.08-0.05)

0.05+0.045

0.095=9.5%

WACC (Book Value)=(35000000/135000000)(0.095)+(100000000/135000000*0.1*0.2)

=0.02463+0.0148=0.03943 =3.943% approximately=4%

WACC(Market Value)=(50000000/15250000)*(0.095)+(102500000/152500000*0.1*0.2)

0.03115+0.01344=0.04459=4.459% approximately=4.46%.